# On Solving Asymmetric Diagonally Dominant Linear Systems in Sublinear Time

### v1

@@ 1.9 Future Directions
Second, establishing stronger lower bounds through unified techniques represents a promising direction. Current spectral gap lower bound for SDD solvers and graph parameter lower bounds for PageRank computation rely on different proofs, establishing lower bounds for the number of probes into $\boldsymbol{b}$ and into the graph structure, respectively. Combining these techniques may yield generalized and tighter lower bounds. Crucially, for local PageRank and effective resistance computation, while the dependence on the decay factor and the spectral gap is crucial [FRS+19, FY22, CCP23, CWW25], surprisingly, no lower bounds on them are known. In this regard, the spectral gap lower bound for SDD solvers may provide inspiration for proving such lower bounds.
@@ 2.1 More Notations and Facts
The pseudoinverse of $\mathbf{M}\in\mathbb{R}^{n\times n}$ satisfies $\mathbf{M}^{+}=\left(\mathbf{M}|_{\ker(\mathbf{M})^{\perp}}\right)^{-1}\P_{\operatorname{range}(\mathbf{M})}$, where $\mathbf{M}|_{\ker(\mathbf{M})^{\perp}}$ is an invertible linear map from $\ker(\mathbf{M})^{\perp}$ to $\operatorname{range}(\mathbf{M})$ and $\P_{\operatorname{range}(\mathbf{M})}$ is the orthogonal projection from $\mathbb{R}^{n}$ onto $\operatorname{range}(\mathbf{M})$ (see, e.g., [Axl24, Section 6C]).
@@ 2.2 ForwardPush and BackwardPush on Graphs
ForwardPush [ACL06, ACL07] and BackwardPush [ABC+07, ABC+08] are two local exploration algorithms for estimating random-walk probabilities on graphs. They are originally developed for computing PPR and PageRank contribution vectors, respectively. [^4] Their original names are ApproximatePageRank and ApproxContributions, respectively. For our purpose, we describe multi-level variants of both algorithms for PPR computation, which are similar to their original forms in spirit, but can be readily adapted to estimate multi-step random-walk probabilities and align better with our more general framework. We refer interested readers to [YWW+24] for a more detailed treatment of the original version of ForwardPush and BackwardPush.
@@ 10    $\boldsymbol{r}^{(\ell)}(v)\leftarrow 0$
Algorithm 2 $\textup{{BackwardPush}}(G,t,\alpha,L,r_{\max})$ [ABC+07, ABC+08]
@@
However, as summarized in [YWW+24], meaningful worst-case bounds on unweighted directed graphs are known only for ForwardPush’s running time and BackwardPush’s approximation error. On unweighted undirected graphs, ForwardPush also provides degree-normalized accuracy guarantees, while BackwardPush enjoys worst-case running time bounds. We do not elaborate on these properties since our unified analysis of the general Push primitive subsumes them.
As noted in [YWW+24], a unified understanding of ForwardPush and BackwardPush remains elusive despite their extensive study and applications. In particular, it is unclear why these algorithms use different conditions for performing the push operations and exhibit different behaviors on directed graphs regarding accuracy and running time bounds. In this work, we provide a unified perspective by viewing both algorithms as instances of the general Push primitive, explaining their different push strategies and properties.
@@ 3 Other Related Work
A vast literature exists on nearly-linear-time Laplacian solvers and their extensions, including solvers for undirected Laplacian/SDD systems (e.g., [ST04, ST14, JS21]) and directed Laplacian/RDD/CDD systems (e.g., [CKP+16, CKK+18, JSS+25]). These solvers achieve nearly-linear time complexity in $\operatorname{nnz}(\mathbf{M})$ with polylogarithmic dependence on $1/\varepsilon$ and the condition number. The development of global RDD/CDD solvers relies on reductions from solving RDD/CDD systems to solving Eulerian Laplacian systems, combined with efficient methods for computing PPR vectors with small $\alpha$ and the stationary distribution of random walk on graphs [CKP+16].
However, the global techniques from algorithmic linear algebra and known reductions to the Eulerian case do not directly apply to the local setting, revealing a fundamental distinction between global and local RDD/CDD solvers. In fact, the lower bounds established in [AKP18] and this work demonstrate that local SDD solvers require polynomial dependence on $1/\varepsilon$ and $1/\gamma$, indicating a separation between global and local SDD solvers.
@@
[DGT17] develops probabilistic logspace solvers for certain classes of directed Laplacian systems. Their method also relies on approximating truncated Neumann series, but they bound the truncation error using spectral radius and Jordan normal form analysis, which yield truncation parameters of at least $n^{2}$. Such a huge truncation parameter makes their algorithm and analysis inapplicable to the sublinear-time setting.
@@
The idea of using random-walk sampling to solve linear systems dates back to the von Neumann-Ulam algorithm for approximating matrix inversion [FL50, Was52]. The bidirectional method for estimating random-walk probabilities on graphs is first proposed in [LBGC14], which is inspired by property testing techniques [GR11, KPS13] and later simplied by the BiPPR framework [LBG16].
@@
The bidirectional idea has been widely applied to compute PageRank [LBG16, BL15, LBG15, BPP18, BPP23, WWY24, WWWY24], effective resistance [CWW25, YLLW25], heat kernel [BPP18, BPP23], and Markov Chain transition probability [BL15]. Among them, [WWWY24] proves that the simple BiPPR framework computes single-node PageRank on unweighted directed graphs in optimal time complexity (in terms of $n$ and $m$). Their analysis relies on a new complexity bound of using BackwardPush to approximate PageRank contributions, which is parameterized by the PageRank value of the target node. Recently, [YLLW25] shows that the bidirectional technique can yield simpler and faster algorithms for constructing effective resistance sketch on expander graphs.
@@
[SBL16] uses the bidirectional method to estimate a single element in the product of a matrix power and a vector, which relates to our approach of estimating $\frac{1}{2}\boldsymbol{t}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$. However, they only derive an average-case complexity bound under some bounded-norm conditions and discuss its applications to solving PSD systems. In contrast, we conduct a more comprehensive study of this problem and apply it to solving RDD/CDD systems.
PageRank and PPR have been extensively studied and widely applied; we refer interested readers to the surveys [LM03, Gle15, YWW+24]. More lower bounds for PageRank computation can be found in [BPP23, YWW+24] and references therein.
@@
Effective resistance is ubiquitous in spectral graph theory [DS84, Lov93, SS11, JSS+25]. A line of work [PLYG21, YT23, CWW25, YLLW25] focuses on locally estimating effective resistance between two nodes through multi-step random-walk probabilities. [CCP23] studies this problem on non-expander graphs and establishes strong lower bounds, though it does not explicitly give a lower bound on the spectral gap of the graph. Recently, [YLLW25] provides a lower bound on the relative error parameter for this problem.
4 Formulation of $\boldsymbol{x}^{\ast}$ and the $p$-Norm Gap
@@
In this section, we prove Theorems 1.1, 1.2, and 1.3 and give further explanations on our formulations of $\boldsymbol{x}^{\ast}$ and the $p$-norm gap.
@@ 5 Random-Walk Sampling
We first focus on RDD systems and then transfer the results to CDD systems by transposing $\mathbf{M}$ and the expression of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$. When $\mathbf{M}$ is RDD, $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}^{\top}\right|\right)$ is row substochastic and we can estimate
@@
by generating random-walk sampling from $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$. See Algorithm 3 for a pseudocode. In each sample, we first sample a walk length $\ell$ from [0,L1][0,L-1] uniformly at random and sample a source coordinate from the distribution $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$. Then we simulate a lazy random walk for $\ell$ steps starting from $v$ according to the transition matrix $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)$. Specifically, at each step from $v$, with probability $\frac{1}{2}$ the walk stays put at $v$, with probability $\frac{|\mathbf{A}_{\mathbf{M}}(u,v)|}{2d_{\mathbf{M}}(v)}$ the walk moves to each $u\in[n]$, and with the remaining probability $\frac{1}{2}-\sum_{u\in[n]}\frac{|\mathbf{A}_{\mathbf{M}}(u,v)|}{2d_{\mathbf{M}}(v)}$ the walk terminates. Note that these probabilities lie in $[0,1]$ since $\mathbf{M}$ is RDD. Additionally, we need to keep track of the product of the signs of the initial entry in $\boldsymbol{t}$ and the entries in $\mathbf{A}_{\mathbf{M}}$ along the walk (where stay-put steps have sign $1$), which is denoted as $\sigma$ in the pseudocode. After $\ell$ steps, if the walk has not terminated and is at coordinate $v$, we take the value $\sigma\cdot\frac{1}{2}\|\boldsymbol{t}\|_{1}\cdot\frac{\boldsymbol{b}(v)}{d_{\mathbf{M}}(v)}\cdot L$ as the estimate of this sample. We repeat this process for $n_{\mathrm{s}}$ independent samples and return the average as the final estimate $\hat{x}$.
@@ 11    $v\leftarrow v^{\prime}$
We first establish the unbiasedness of the sampling.
@@
We use Algorithm 3. The algorithm returns $\hat{x}$ as the average of $n_{\mathrm{s}}$ independent samples, where each sampled value has absolute value at most $\frac{1}{2}\|\boldsymbol{t}\|_{1}\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}L$. Therefore, by Lemma 5.1 and the Hoeffding bound (Theorem A.1), $\Pr\left\{\left|\hat{x}-\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}\right|\geq\frac{1}{2}\varepsilon\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\right\}$ is upper bounded by
@@
As each sample simulates at most $L$ steps of the random walk, and each step takes $O\big{(}f_{\mathrm{row}}(\mathbf{M})\big{)}$ time, the time complexity is $O\big{(}f_{\mathrm{row}}(\mathbf{M})L\cdot n_{\mathrm{s}}\big{)}=O\left(f_{\mathrm{row}}(\mathbf{M})\|\boldsymbol{t}\|_{1}^{2}L^{3}/\varepsilon^{2}\right)$, as desired. ∎
@@
On the other hand, if $\mathbf{M}$ is CDD, then $\mathbf{M}^{\top}$ is RDD and $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}\right)$ is row substochastic. As we can transpose the expression of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$ to obtain
@@ 7    $\boldsymbol{r}^{(\ell)}(v)\leftarrow 0$
We prove the lemma by induction on the number of push operations performed. As the base case, initially only $\boldsymbol{r}^{(0)}=\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ contains nonzero entries, so the right-hand side of the invariant equation equals $\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$, which equals the left-hand side of the equation, as desired. For the inductive step, we examine the alteration of the right-hand side of the equation incurred by a single push operation on $v\in[n]$ at level $\ell^{\prime}$, and it suffices to check that the alteration is $\boldsymbol{0}$. The first step changes $\boldsymbol{p}^{(\ell^{\prime})}(v)$ from 0 to $\boldsymbol{r}^{(\ell^{\prime})}(v)$, so $\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}$ is increased by $\boldsymbol{r}^{(\ell^{\prime})}(v)\boldsymbol{e}_{v}$. The second step increases $\boldsymbol{r}^{(\ell^{\prime}+1)}$ by $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\left(\boldsymbol{r}^{(\ell^{\prime})}(v)\boldsymbol{e}_{v}\right)$, which in term increases the right-hand side of the invariant equation by $\boldsymbol{r}^{(\ell)}(v)$ times
@@
Observe that each time a push operation is performed on a coordinate $v\in[n]$, the value of $\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left|\boldsymbol{p}^{(\ell)}\right|$ is increased by at least $r_{\max}$. However, by Lemma 6.3, $\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left|\boldsymbol{p}^{(\ell)}\right|$ is always upper bounded by $\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}^{\top}\right|\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}|\boldsymbol{b}|$. Therefore, the total number of push operations performed on $v$ is at most $\frac{1}{r_{\max}}\cdot\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}^{\top}\right|\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}|\boldsymbol{b}|$. Since each push operation on $v$ takes $O\left(\big{\|}\mathbf{M}(\cdot,v)\big{\|}_{0}\right)$ time, the lemma follows by summing the cost of the push operations over all $v\in[n]$ and adding the $O\big{(}\|\boldsymbol{b}\|_{0}\big{)}$ time for the initialization. ∎
@@
We run Algorithm 4 with $r_{\max}:=\varepsilon/L^{2}$ and use $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ as the result. Since $\mathbf{M}$ is RDD, by Lemma 6.2, the absolute error between the estimate and $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$ is upper bounded by $\frac{1}{2}\|\boldsymbol{t}\|_{1}L^{2}\cdot r_{\max}\leq\frac{1}{2}\varepsilon\|\boldsymbol{t}\|_{1}$. Since $\mathbf{M}$ is CDD and its nonzero entries have absolute values of $\Omega(1)$, by Lemma 6.5, the time complexity of the Push algorithm is $O\big{(}\|\boldsymbol{b}\|_{0}+\|\boldsymbol{b}\|_{1}L/r_{\max}\big{)}=O\big{(}\|\boldsymbol{b}\|_{0}+\|\boldsymbol{b}\|_{1}L^{3}/\varepsilon\big{)}$. This finishes the proof. ∎
@@
We run Algorithm 4 with $r_{\max}:=\varepsilon\big{/}\big{(}d_{\mathbf{M}}(w)L^{2}\big{)}$. By Lemma 6.2, the absolute error is upper bounded by $\frac{1}{2}\|\boldsymbol{t}\|_{1}L^{2}\cdot r_{\max}=\frac{1}{2}\varepsilon\|\boldsymbol{t}\|_{1}/d_{\mathbf{M}}(w)=\frac{1}{2}\varepsilon\|\boldsymbol{t}\|_{1}\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}$. Using Lemma 6.4, we can upper bound the average push cost over $w$ by

### v2

@@ 1.9 Future Directions
Second, establishing stronger lower bounds through unified techniques represents a promising direction. Current spectral gap lower bound for SDD solvers and graph parameter lower bounds for PageRank computation rely on different proofs, establishing lower bounds for the number of probes into $\boldsymbol{b}$ and into the graph structure, respectively. Combining these techniques may yield generalized and tighter lower bounds. Notably, for local PageRank and effective resistance computation, the dependence on the decay factor and the spectral gap is crucial [FRSCM19, FY22, CCP23, CWW25]. However, surprisingly, no lower bounds on these parameters are known to the best of our knowledge. In this regard, the spectral gap lower bound for SDD solvers may provide inspiration for proving such lower bounds.
@@ 2.1 More Notations and Facts
The pseudoinverse of $\mathbf{M}\in\mathbb{R}^{n\times n}$ satisfies $\mathbf{M}^{+}=\left(\mathbf{M}|_{\ker(\mathbf{M})^{\perp}}\right)^{-1}\mathbf{P}_{\operatorname{range}(\mathbf{M})}$, where $\mathbf{M}|_{\ker(\mathbf{M})^{\perp}}$ is an invertible linear map from $\ker(\mathbf{M})^{\perp}$ to $\operatorname{range}(\mathbf{M})$ and $\mathbf{P}_{\operatorname{range}(\mathbf{M})}$ is the orthogonal projection from $\mathbb{R}^{n}$ onto $\operatorname{range}(\mathbf{M})$ (see, e.g., [Axl24, Section 6C]).
@@ 2.2 ForwardPush and BackwardPush on Graphs
ForwardPush [ACL06, ACL07] and BackwardPush [ABCHMT07, ABCHMT08] are two local exploration algorithms for estimating random-walk probabilities on graphs. They are originally developed for computing PPR and PageRank contribution vectors, respectively. [^4] Their original names are ApproximatePageRank and ApproxContributions, respectively. For our purpose, we describe multi-level variants of both algorithms for PPR computation, which are similar to their original forms in spirit, but can be readily adapted to estimate multi-step random-walk probabilities and align better with our more general framework. We refer interested readers to [YWWWW24] for a more detailed treatment of the original version of ForwardPush and BackwardPush.
@@ 10    $\boldsymbol{r}^{(\ell)}(v)\leftarrow 0$
Algorithm 2 $\textup{{BackwardPush}}(G,t,\alpha,L,r_{\max})$ [ABCHMT07, ABCHMT08]
@@
However, as summarized in [YWWWW24], meaningful worst-case bounds on unweighted directed graphs are known only for ForwardPush’s running time and BackwardPush’s approximation error. On unweighted undirected graphs, ForwardPush also provides degree-normalized accuracy guarantees, while BackwardPush enjoys worst-case running time bounds. We do not elaborate on these properties since our unified analysis of the general Push primitive subsumes them.
As noted in [YWWWW24], a unified understanding of ForwardPush and BackwardPush remains elusive despite their extensive study and applications. In particular, it is unclear why these algorithms use different conditions for performing the push operations and exhibit different behaviors on directed graphs regarding accuracy and running time bounds. In this work, we provide a unified perspective by viewing both algorithms as instances of the general Push primitive, explaining their different push strategies and properties.
@@ 3 Other Related Work
A vast literature exists on nearly-linear-time Laplacian solvers and their extensions, including solvers for undirected Laplacian/SDD systems (e.g., [ST04, ST14, JS21]) and directed Laplacian/RDD/CDD systems (e.g., [CKPPSV16, CKKPPRS18, JSSTZ25]). These solvers achieve nearly linear time complexity in $\operatorname{nnz}(\mathbf{M})$ with polylogarithmic dependence on $1/\varepsilon$ and the condition number. The development of global RDD/CDD solvers relies on reductions from solving RDD/CDD systems to solving Eulerian Laplacian systems, combined with efficient methods for computing PPR vectors with small $\alpha$ and the stationary distribution of random walks on graphs [CKPPSV16]. However, the global techniques from algorithmic linear algebra and known reductions to the Eulerian case do not directly apply to the local setting, revealing a fundamental distinction between global and local RDD/CDD solvers. In fact, the lower bounds established in [AKP18] and this work demonstrate that local SDD solvers require polynomial dependence on $1/\varepsilon$ and $1/\gamma$, indicating a separation between global and local SDD solvers.
@@
[DGT17] develops probabilistic logspace solvers for certain classes of directed Laplacian systems. Their method also relies on approximating truncated Neumann series, but they bound the truncation error using spectral radius and Jordan normal form, which yield truncation parameters of at least $n^{2}$. Such a huge truncation parameter makes their algorithm and analysis inapplicable to the sublinear-time setting. As an aside, in the quantum regime, [Ta-13] presents algorithms for inverting well-conditioned matrices in quantum logspace, but their approaches are not directly applicable to our classical sublinear-time framework.
@@
The idea of using random-walk sampling to solve linear systems dates back to the von Neumann-Ulam algorithm for approximating matrix inversion [FL50, Was52]. The bidirectional method for estimating random-walk probabilities on graphs is first proposed in [LBGC14], which is inspired by property testing techniques [GR11, KPS13] and later simplified by the BiPPR framework [LBG16].
@@
The bidirectional idea has been widely applied to compute PageRank [LBG15, LBG16, BPP18, BPP23, WWY24, WWWY24, BJTWY25, TWWY26], effective resistance [CWW25, YLLW25], heat kernel [BPP18, BPP23], and Markov Chain transition probability [BL15]. Among them, [WWWY24] proves that the simple BiPPR framework computes single-node PageRank on unweighted directed graphs in optimal time complexity (in terms of $n$ and $m$). Their analysis relies on a new complexity bound of BackwardPush, which is parameterized by the PageRank value of the target node. Recently, [YLLW25] shows that the bidirectional technique can yield faster algorithms for constructing effective resistance sketch (as defined in [DKS23]) on expander graphs, and [TWWY26] combines random-walk sampling with a novel randomized local push technique to improve the complexity of estimating single-node PageRank on directed graphs with bounded in-degree.
@@
[SBL16] uses the bidirectional method to estimate a single element in the product of a matrix power and a vector, which relates to our estimation of $\frac{1}{2}\boldsymbol{t}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$. However, they only derive an average-case complexity bound under some bounded-norm conditions and discuss its applications to solving PSD systems. In contrast, we conduct a more comprehensive study of this problem and apply it to solving RDD/CDD systems.
PageRank and PPR have been extensively studied and widely applied; we refer interested readers to the surveys [LM03, Gle15, YWWWW24]. More lower bounds for PageRank computation can be found in [BPP23, YWWWW24, BJTWY25] and references therein. The recent work [BJTWY25] conducts a comprehensive study of various types of PPR estimation problems using different graph access queries under both worst-case and average-case settings. For constant decay factors, they provide nearly tight complexity upper and lower bounds for achieving constant relative error guarantees when the target value is above a given threshold.
@@
Effective resistance is ubiquitous in spectral graph theory [DS84, Lov93, SS11, JSSTZ25]. A line of work [PLYG21, YT23, CWW25, YLLW25] focuses on locally estimating single-pair effective resistance through multi-step random-walk probabilities. [CCP23] studies this problem on non-expander graphs and establishes strong complexity lower bounds, though it does not explicitly give a lower bound on the spectral gap of the graph. Recently, [YLLW25] provides a complexity lower bound on the relative error parameter for this problem. Besides, a line of work [LS23, DKS23, YLLW25] studies the construction of effective resistance sketches. To this end, [LS23] leverages random-walk sampling, [DKS23] uses count sketches and SDD solvers, and [YLLW25] employs a bidirectional approach.
4 Formulation of $\boldsymbol{x}^{\ast}$ and the $p$-Norm Gaps
@@
In this section, we prove Theorems 1.1, 1.2, and 1.3 and give further explanations on our formulations of $\boldsymbol{x}^{\ast}$ and the $p$-norm gaps.
@@ 5 Random-Walk Sampling
We first focus on RDD systems and transfer the results to CDD systems by transposing $\mathbf{M}$ and the expression of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$ at the end of this section. When $\mathbf{M}$ is RDD, $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}^{\top}\right|\right)$ is row substochastic and we can estimate
@@
by generating random-walk sampling from $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$. See Algorithm 3 for a pseudocode. In each sample, we first sample a walk length $\ell$ from [0,L1][0,L-1] uniformly at random and sample a source coordinate $v$ from the distribution $|\boldsymbol{t}|/\|\boldsymbol{t}\|_{1}$. Then we simulate a lazy random walk for $\ell$ steps starting from $v$ according to the transition matrix $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}^{\top}\right|\right)$. Specifically, at each step from $v$, with probability $\frac{1}{2}$ the walk stays put at $v$, with probability $\frac{|\mathbf{A}_{\mathbf{M}}(u,v)|}{2d_{\mathbf{M}}(v)}$ the walk moves to each $u\in[n]$, and with the remaining probability $\frac{1}{2}-\sum_{u\in[n]}\frac{|\mathbf{A}_{\mathbf{M}}(u,v)|}{2d_{\mathbf{M}}(v)}$ the walk terminates. Note that these probabilities lie in $[0,1]$ since $\mathbf{M}$ is RDD. Additionally, we keep track of the product of the signs of the initial entry in $\boldsymbol{t}$ and the entries in $\mathbf{A}_{\mathbf{M}}$ along the walk (where stay-put steps have sign $1$), which is denoted as $\sigma$ in the pseudocode. After $\ell$ steps, if the walk has not terminated and is at coordinate $v$, we take the value $\sigma\cdot\frac{1}{2}\|\boldsymbol{t}\|_{1}\cdot\frac{\boldsymbol{b}(v)}{d_{\mathbf{M}}(v)}\cdot L$ as the estimate of this sample. We repeat this process for $n_{\mathrm{s}}$ independent samples and return the average as the final estimate $\hat{x}$.
@@ 11    $v\leftarrow v^{\prime}$
We first establish the unbiasedness of the sampling scheme.
@@
We use random-walk sampling as in Algorithm 3. The algorithm returns $\hat{x}$ as the average of $n_{\mathrm{s}}$ independent samples, where each sampled value has absolute value at most $\frac{1}{2}\|\boldsymbol{t}\|_{1}\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}L$. Therefore, by Lemma 5.1 and the Hoeffding bound (Theorem A.1), $\Pr\left\{\left|\hat{x}-\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}\right|\geq\frac{1}{2}\varepsilon\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}\right\}$ is upper bounded by
@@
As each sample simulates at most $L$ steps of random walk, and each step takes $O\big(f_{\mathrm{row}}(\mathbf{M})\big)$ time, the time complexity is $O\big(f_{\mathrm{row}}(\mathbf{M})L\cdot n_{\mathrm{s}}\big)=O\left(f_{\mathrm{row}}(\mathbf{M})\|\boldsymbol{t}\|_{1}^{2}L^{3}/\varepsilon^{2}\right)$, as desired. ∎
@@
On the other hand, if $\mathbf{M}$ is CDD, then $\mathbf{M}^{\top}$ is RDD and $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}\right|\right)$ is row substochastic. As we can transpose the expression of $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$ to obtain
@@ 7    $\boldsymbol{r}^{(\ell)}(v)\leftarrow 0$
We prove the lemma by induction on the number of push operations performed. As the base case, initially only $\boldsymbol{r}^{(0)}=\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$ contains nonzero entries, so the right-hand side of the invariant equation equals $\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}$, which equals the left-hand side of the equation, as desired. For the inductive step, we examine the alteration of the right-hand side of the equation incurred by a single push operation on $v\in[n]$ at level $\ell^{\prime}$, and it suffices to check that the alteration is $\boldsymbol{0}$. The first step changes $\boldsymbol{p}^{(\ell^{\prime})}(v)$ from 0 to $\boldsymbol{r}^{(\ell^{\prime})}(v)$, so $\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}$ is increased by $\boldsymbol{r}^{(\ell^{\prime})}(v)\boldsymbol{e}_{v}$. The second step increases $\boldsymbol{r}^{(\ell^{\prime}+1)}$ by $\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\mathbf{A}_{\mathbf{M}}^{\top}\right)\left(\boldsymbol{r}^{(\ell^{\prime})}(v)\boldsymbol{e}_{v}\right)$, which in turn increases the right-hand side of the invariant equation by $\boldsymbol{r}^{(\ell)}(v)$ times
@@
Observe that each time a push operation is performed on a coordinate $v\in[n]$, the value of $\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left|\boldsymbol{p}^{(\ell)}\right|$ is increased by at least $r_{\max}$. However, by Lemma 6.3, $\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left|\boldsymbol{p}^{(\ell)}\right|$ is always upper bounded by $\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}^{\top}\right|\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}|\boldsymbol{b}|$. Therefore, the total number of push operations performed on $v$ is at most $\frac{1}{r_{\max}}\cdot\boldsymbol{e}_{v}^{\top}\sum_{\ell=0}^{L-1}\left(\frac{1}{2}\left(\mathbf{I}+\mathbf{D}_{\mathbf{M}}^{-1}\left|\mathbf{A}_{\mathbf{M}}^{\top}\right|\right)\right)^{\ell}\mathbf{D}_{\mathbf{M}}^{-1}|\boldsymbol{b}|$. Since each push operation on $v$ takes $O\left(\big\|\mathbf{M}(\cdot,v)\big\|_{0}\right)$ time, the lemma follows by summing the cost of the push operations over all $v\in[n]$ and adding the $O\big(\|\boldsymbol{b}\|_{0}\big)$ time for initialization. ∎
@@
We run Push with $r_{\max}:=\varepsilon/L^{2}$ and use $\frac{1}{2}\boldsymbol{t}^{\top}\left(\sum_{\ell=0}^{L-1}\boldsymbol{p}^{(\ell)}+\boldsymbol{r}^{(L-1)}\right)$ as the result. Since $\mathbf{M}$ is RDD, by Lemma 6.2, the absolute error between the estimate and $\boldsymbol{t}^{\top}\boldsymbol{x}^{\ast}_{L}$ is upper bounded by $\frac{1}{2}\|\boldsymbol{t}\|_{1}L^{2}\cdot r_{\max}\leq\frac{1}{2}\varepsilon\|\boldsymbol{t}\|_{1}$. Since $\mathbf{M}$ is CDD and its nonzero entries have absolute values of $\Omega(1)$, by Lemma 6.5, the time complexity of the Push algorithm is $O\big(\|\boldsymbol{b}\|_{0}+\|\boldsymbol{b}\|_{1}L/r_{\max}\big)=O\big(\|\boldsymbol{b}\|_{0}+\|\boldsymbol{b}\|_{1}L^{3}/\varepsilon\big)$. This finishes the proof. ∎
@@
We run Push with $r_{\max}:=\varepsilon\big/\big(d_{\mathbf{M}}(w)L^{2}\big)$. By Lemma 6.2, the absolute error is upper bounded by $\frac{1}{2}\|\boldsymbol{t}\|_{1}L^{2}\cdot r_{\max}=\frac{1}{2}\varepsilon\|\boldsymbol{t}\|_{1}/d_{\mathbf{M}}(w)=\frac{1}{2}\varepsilon\|\boldsymbol{t}\|_{1}\left\|\mathbf{D}_{\mathbf{M}}^{-1}\boldsymbol{b}\right\|_{\infty}$. Using Lemma 6.4, we can upper bound the average push cost over $w$ by